# The Diophantine problem in isotropic reductive groups

### v1

@@ 5 Main results
    for $x,y\in A$, where $A$ is a composition algebra of rank $8$ or an Azumaya algebra over $K$. Hence $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)=U_{\mathrm{e}_{1}-\mathrm{e}_{3}}(x)$ (i.e. all maps $U_{\alpha}$ coincide) and $U_{\alpha}(x)=x\,U_{\alpha}(1)=U_{\alpha}(1)\,x$, so $U_{\alpha}(1)$ lies in the center of $A$, and this center is precisely $K$.
  • In the orthogonal-like case ($\mathsf{B}_{n,2}$, ${{}^{1}\!{\mathsf{D}_{n,2}^{(d)}}}$, ${{}^{2}\!{\mathsf{D}_{n,2}^{(d)}}}$) we have
    $U_{\mathrm{e}_{2}}(m)\,r=m\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)=U_{\mathrm{e}_{1}}(mr)$
    for $r\in R$ and $m\in M$, where $R$ is an Azumaya algebra over $K$ and $M_{R}$ is a faithfully projective module. We have $U_{\mathrm{e}_{1}}(m)=U_{\mathrm{e}_{2}}(m)=m\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)$ and $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)=r\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)=U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,r$, so $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in K=\operatorname{C}(R)$.
  • In the remaining symplectic-like case we have
    $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)\,u=r\,U_{2\mathrm{e}_{2}}(u)=U_{\mathrm{e}_{1}+\mathrm{e}_{2}}(ru)$
    for $r\in R$ and $u\in\Lambda$, where $R$ is an Azumaya algebra over $K$ or its quadratic étale extension and $\Lambda\leq R$ is a form parameter containing invertible element $\iota$. It follows that $U_{2\mathrm{e}_{2}}(u)=U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,u$, $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)=r\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)$, and $U_{\mathrm{e}_{1}+\mathrm{e}_{2}}(r)=r\iota^{-1}\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,\iota$. Then $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,u\iota^{-1}=u\iota^{-1}\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)$ and $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\Lambda\iota^{-1}$. In the subcase ${{}^{2}\!{\mathsf{A}_{n,2}^{(d)}}}$ the group $\Lambda=R_{0}$ is an Azumaya algebra over $K$ (so $R=R_{0}\otimes\operatorname{C}(R)$) and $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\Lambda\iota^{-1}\cap\operatorname{C}(R)=K$. In the subcase $\mathsf{C}_{n,2}^{(d)}$ the form parameter is
    $\Lambda=\{x\in\mathrm{M}(d,K)\mid x_{ij}=x_{d+1-j,d+1-i}\}$
    and $\iota=1$tale locally, so $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\operatorname{C}_{R}(\Lambda)=K$. In the last subcase ${{}^{1}\!{\mathsf{D}_{n,2}^{(d)}}}$ we have
    $\Lambda=\{x\in\mathrm{M}(d,K)\mid x_{ij}=-x_{d+1-j,d+1-i},x_{i,d+1-i}=0\}$
    and $\iota=e_{11}+\ldots+e_{\frac{d}{2},\frac{d}{2}}-e_{\frac{d}{2}+1,\frac{d}{2}+1}-\ldots-e_{dd}$tale locally, so again $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\operatorname{C}_{R}(\Lambda\iota^{-1})\cap\Lambda\iota^{-1}=K$ (if $d\geq 2$, then already $\operatorname{C}_{R}(\Lambda\iota^{-1})=K$).
@@
It remains to define ring operations on $\widetilde{K}$. The group operation is directly induced from $G(K)$. Finally, the product of $(k_{\alpha})_{\alpha}$ and $(l_{\alpha})_{\alpha}$ is such a family $(m_{\alpha})_{\alpha}$ that
$[k_{\alpha}(x_{\alpha}),l_{\beta}(x_{\beta})]\equiv m_{\alpha+\beta}([x_{\alpha},x_{\beta}])\pmod{\operatorname{C}(G)(K)\,\prod_{\begin{subarray}{c}i\alpha+j\beta\in\Phi\\ i,j\geq 1\\ (i,j)\neq(1,1)\end{subarray}}G_{i\alpha+j\beta}}$
@@
for $x_{\alpha}\in X_{\alpha}$, $x_{\beta}\in X_{\beta}$, and roots $\alpha$, $\beta$ such that $\alpha+\beta$ is also a root. It suffices to impose this condition only for one pair $(\alpha,\beta)$. ∎
@@
Let $K$ be a commutative ring with an enumeration by $\mathbb{N}$ or its finite subset such that the ring operations are computable. Let also $G$ be an isotropic reductive group over $K$ such that the rank of $\Phi$ is at least $2$. Then the Diophantine problems for $K$ and $G(K)$ are equivalent, i.e. they reduce to each other by algorithms. Here the enumeration on $G(K)$ is obtained from some closed embedding $G\subseteq\mathbb{A}^{n}_{K}$, it is independent of the embedding up to a computable permutation.
@@
Clearly, the group operations on $G(K)$ are also computable. Every Diophantine subset of $G(K)^{k}$ may be considered as a Diophantine subset of $K^{kn}$ using the e-interpretation of $G(K)$ in $K$ and this transformation is clearly computable in terms of defining formulae. In the other direction apply theorem 5. ∎

### v2

@@ 5 Main results
    for $x,y\in\operatorname{M}(d,K)$. Hence $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)=k_{\mathrm{e}_{1}-\mathrm{e}_{3}}(x)=k_{\mathrm{e}_{2}-\mathrm{e}_{3}}(x)$ (i.e. all maps $k_{\alpha}$ coincide) and $k_{\alpha}(x)=x\,k_{\alpha}(1)=k_{\alpha}(1)\,x$, so $k_{\alpha}(1)$ lies in the center of $\operatorname{M}(d,K)$, and this center is precisely $K$.
  • The case $^{1}\!{\mathsf{E}_{6,2}^{28}}$} is similar to the previous one. We have
    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=k_{\mathrm{e}_{1}-\mathrm{e}_{3}}(xy)=x\,k_{\mathrm{e}_{2}-\mathrm{e}_{3}}(y)$
    for $x,y\in\mathrm{Z}(K)$, where $\mathrm{Z}(K)$ is the split octonion algebra over $K$. Hence all maps $k_{\alpha}$ coincide and $k_{\alpha}(x)=x\,k_{\alpha}(1)=k_{\alpha}(1)\,x$, so $k_{\alpha}(1)\in K$.
  • In the case $^{2}\!{\mathsf{A}_{4d-1,2}^{(d)}}$} we have
    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x,y)\circ z=(x,y)\circ k_{2\mathrm{e}_{2}}(z)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}((x,y)\circ z)$
    for $x,y,z\in\operatorname{M}(d,K)$, where $(x,y)\circ z=(zx,yz)$. It follows that $k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(x,y)=(x,y)\circ k_{\mathrm{e}_{2}}(z)$ and $k_{\mathrm{e}_{2}}(z)=zk_{\mathrm{e}_{2}}(1)=k_{\mathrm{e}_{2}}(1)z$, so $k_{\mathrm{e}_{2}}(1)\in K$.
  • In the case $\mathsf{C}_{2d,2}^{(d)}$ we have
    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=x\,k_{2\mathrm{e}_{2}}(y)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(xy)$
    for $x\in\operatorname{M}(d,K)$ and symmetric matrix $y=y^{\mathrm{t}}\in\operatorname{M}(d,K)$. Then $k_{2\mathrm{e}_{2}}(y)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(y)$ and $k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(x)=k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)=x\,k_{2\mathrm{e}_{2}}(1)$. So $k_{2\mathrm{e}_{2}}(y)=k_{2\mathrm{e}_{2}}(1)\,y=y\,k_{2\mathrm{e}_{2}}(y)$, and it easily follows that $k_{2\mathrm{e}_{2}}(1)\in K$.
  • In the last case $^{1}\!{\mathsf{D}_{2d,2}^{(d)}}$} we have
    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=x\,k_{2\mathrm{e}_{2}}(y)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(xy)$
    for $x\in\operatorname{M}(d,K)$ and alternating $y\in\operatorname{M}(d,K)$ (i.e. $y_{ii}=0$ and $y_{ij}=-y_{ji}$). Since $d\geq 2$ is even (actually, a power of $2$), there is an invertible alternating matrix $J\in\operatorname{M}(d,K)$, say, the matrix of the split symplectic form. As in the previous case, $k_{2\mathrm{e}_{2}}(y)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(y)$, $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)=x\,k_{2\mathrm{e}_{2}}(J)\,J^{-1}$, $k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(x)=xJ^{-1}\,k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(J)$. Then $k_{2\mathrm{e}_{2}}(y)=k_{2\mathrm{e}_{2}}(J)\,J^{-1}y=yJ^{-1}\,k_{2\mathrm{e}_{2}}(J)$ and $k_{2\mathrm{e}_{2}}(J)$ is alternating, so it easily follows that $k_{2\mathrm{e}_{2}}(J)\,J^{-1}\in K$.
@@
It remains to define ring operations on $\widetilde{K}$. The group operation is directly induced from $\mathfrak{g}_{\alpha}$. The product of $(k_{\alpha})_{\alpha}$ and $(l_{\alpha})_{\alpha}$ is such a family $(m_{\alpha})_{\alpha}$ that
$k_{\alpha}(x_{\alpha}),l_{\beta}(x_{\beta})]=m_{\alpha+\beta}([x_{\alpha},x_{\beta}])$
@@
for $x_{\alpha}\in X_{\alpha}$, $x_{\beta}\in X_{\beta}$, and roots $\alpha$, $\beta$ such that $\alpha+\beta$ is also a root. It suffices to impose this condition only for one pair $(\alpha,\beta)$ with long $\alpha$. ∎
@@
Let $K$ be a commutative ring with an enumeration by $\mathbb{N}$ or its finite subset such that the ring operations are computable. Let also $G$ be an isotropic reductive group scheme over $K$ such that the rank of $\Phi$ is at least $2$. Then the Diophantine problems for $K$ and $G(K)$ are equivalent, i.e. they reduce to each other by algorithms. Here the enumeration on $G(K)$ is obtained from some closed embedding $G\subseteq\mathbb{A}^{n}_{K}$, it is independent of embedding up to a computable permutation.
@@
Clearly, the group operations on $G(K)$ are also computable. Every Diophantine subset of $G(K)^{k}$ can be considered as a Diophantine subset of $K^{kn}$ using the e-interpretation of $G(K)$ in $K$ and this transformation is clearly computable in terms of defining formulae. In the other direction apply theorem 5. ∎